# The Effect of Predation on the Dynamics of Chronic Wasting Disease in Deer

### v1

@@ 1 Abstract
Chronic Wasting Disease (CWD) is a neurological disease impacting deer, elk, moose, and other cervid populations and is caused by a misfolded protein known as a prion. CWD is difficult to control due to the persistence of prions in the environment. Prions can remain infectious for more than a decade and have been found in soil as well as other environmental vectors, such as ticks and plants. Here, we provide a bifurcation analysis of a mathematical model of CWD spread in a cervid population, and use a modification of the Gillespie algorithm to explore if wolves can be used as an ecological control strategy to limit the spread of the disease in several relevant scenarios. We then analytically compute the probability that the disease spreads given one infected member enters a fully healthy population and the probability of elimination, given a fully susceptible population and remaining prions in the environment. From our analysis, we conclude that wolves can be used as an effective control strategy to limit the spread of CWD in cervid populations, and hunting or other means of lowering the susceptible population are beneficial to controlling the spread of CWD, although it is important to note that inferring biologically relevant parameters from the existing data is an ongoing challenge for this system.
@@ 2 Introduction
Chronic Wasting Disease (CWD) is a prion-mediated disease that has been observed in cervids, such as elk, deer, and moose. Prions are misfolded proteins that can cause neurological disease, such as CWD, but also Creutzfeldt-Jakob disease in humans. CWD is transmitted both directly between susceptible animals and indirectly through environmental exposure. Cervids can become infected through contact with saliva, blood, and other bodily fluids of an infected animal or through indirect environmental infections from water sources, plants, soil, or potentially vector-borne insect infections. In recent years, wild populations of cervids have exhibited high levels of CWD prevalence. The 2024 Chronic Wasting Disease Surveillance Report from the Wyoming Game and Fish Department reported that the prevalence of CWD in some populations of mule deer is as high as 66.3%  [^1] The report can be found here https://wgfd.wyo.gov/wyoming-wildlife/wildlife-disease-and-health/chronic-wasting-disease using the drop-down Report menu at the bottom of the website.. CWD has proven difficult to control or eradicate, due to the time scale associated with the persistence of CWD prions in the environment. Prions decay slowly and can remain latent and infectious for more than a decade [8, 31]. Many mathematical models have been developed to help understand the disease spread in deer populations, evaluate the efficacy of control strategies, and infer the modes of transmission. We briefly review several of the published models, but our summary should not be considered exhaustive.
@@
One of the earlier discrete population models examined the spread of CWD and control strategies [21, 11], but the modeling assumptions of the models, in particular frequency-dependent transmission, were reevaluated in the follow-up paper [26]. Miller and co-authors developed and attempted parameter estimation of competing SIR-type models that describe transmission of CWD in a population of cervids through different mechanisms—though the models appear to exhibit a lack of identifiability, which we briefly comment on in the Discussion [20]. They found support for indirect transmission of CWD using a model selection approach [20]. The model presented in [20] was further modified for further mathematical and control-based investigations.
@@
Sharp and Pastor modified the model presented in [20] to include logistic growth in the susceptible population and performed a mathematical analysis of the system [27]. Wild and co-authors later extended the model presented in [20] to include the effects of predation from wolves [32] and analyzed the system in a non-oscillatory parameter regime, using parameter values estimated in [20] and other studies. An earlier report to the National Park Service by Hobbs also analyzed wolf-predation as a control strategy using a similar model system in an oscillatory parameter regime [12]. Barbera and Pollino performed a mathematical analysis of a PDE [4] that was inspired by the model presented in [27]. Reyes and co-workers designed a spatiotemporal causal inference scheme to examine the effects of culling strategies using a PDE-based approach [18]. Potapov and co-authors used a complex age-structure differential equation model to examine the influence of harvesting strategies on population dynamics and disease suppression [24]. Vasilyva and co-authors performed a mathematical analysis of a differential equation model that describes CWD spread in a population of deer, including both indirect and direct transmission [30]. They also attempted a parameter estimation of the system, using upper and lower bounds for parameter estimation based on estimates of parameters found in [20]. A different study examined the influence of culling strategies to control CWD spread using a differential equation model that incorporated seasonality into a population model of cervids [23]. Others created an integrodifference equation model that incorporated direct, indirect, and long-distance dispersal from juvenile deer to understand CWD spread in a population of white-tailed deer [19]. A different group looked at harvest strategies in a mathematical model that split cervids by sex [3, 2]. Control investigations into CWD transmission in deer populations are summarized in [29].
@@ 4 Stochastic simulation
Here, $E$ remains a continuous, not discrete, variable, and so is governed by the differential equation $\frac{dE}{dt}=\frac{\epsilon A_{0}}{A}i-\mu_{e}E$. We would like to use these reaction rates with the Gillespie algorithm [10] to do stochastic simulations. However, since $E$ is continuously changing in time, the Gillespie algorithm is not directly applicable. Instead, we use “Poisson thinning” to account for this fact, as follows[16][22][25]: Since between reactions, $E(t)$ is a monotone function, bounded between $E(t_{0})$ and $\epsilon\frac{iA_{0}}{\mu_{e}A}$, where $t_{0}$ is the last reaction time, we let $E^{*}={\rm max}[E(t_{0}),\epsilon\frac{iA_{0}}{\mu_{e}A}]$, and introduce a modified reaction $R_{3}^{*}$ (since $R_{3}$ depends on $E$), with reaction rate $r_{3}^{*}=\gamma sE^{*}$ (which is the fastest possible reaction rate), and then proceed with the usual Gillespie determination of next reaction time and next reaction. However, whenever the next reaction is to be $R_{3}^{*}$ at time $t^{*}$, the reaction is split into two, only one of which is implemented: a null reaction for which no change to $s$ or $i$ is made with probability $\frac{E^{*}-E(t^{*})}{E^{*}}$, and the reaction that reduces $s$ by one and increases $i$ by one with probability $\frac{E(t^{*})}{E^{*}}$.
@@ 4.3 Probability of Elimination
is the probability that there will not be another outbreak. Obviously, the probability that there will not be another outbreak goes to zero as ${\gamma_{e}\over\mu_{e}}sE_{0}$ becomes large. This once again shows the risk associated with overpopulation of susceptible deer (large $s$), and the benefit of having wolves to help control the population size of healthy deer, even when there are no infected animals present.
@@ 5 Discussion
The prions responsible for CWD are known to exist in soil, and can be measured from soil samples [15]. In practice, estimating $E$ through soil samples would likely be expensive, infeasible on a large scale, and miss some of the sources of the prions in the environment. The prions responsible for CWD can be found in multiple environmental reservoirs beyond soil. It was recently shown that ticks can harbor levels of prions that could cause transmission of CWD [13]. Prions were also recently shown to exist in plants at relevant levels for transmission of CWD [5]. For these reasons, getting estimates of the prion density will be difficult. A different approach to ensure that unique parameters can be found is to reparameterize the model to remove the structural identifiability issue ahead of parameter estimation, but this can alter the interpretability of state variables of the model or rate parameters. Of course, whether unique parameter estimates are needed depends on the how the model will be used.
@@
We chose to use an ODE model and did not consider the role of possible spatial inhomogeneity. However, the role of space may be important in understanding the spread of CWD in deer. Currently, little is known about the spatial distribution of prions. One would imagine the spatial distribution of the prion is somewhat “patchy” in space because an infected animal carcass contains a high concentration, which would then diffuse in the local environment. However, some degree of spatial mixing is likely to occur. For example, rivers could be an agent by which the prions could spread in the environment. It is also known that the animals eating the infected carcass, such as wolves, ravens, and other scavengers, will spread the prions throughout the environment through feces. In our model, we took the term $\xi W$ in Eq. 9 to be zero throughout our analysis, but taking this parameter to be non-zero would begin to capture prion spread by wolves. Currently, less is known about the spread of prions through insects and plants, or if infected plants can pass the prions inter-generationally through seeds, but these are other theoretical mechanisms by which prions could become well-mixed in the environment. More research is needed to understand the spatial distribution of prions in the environment and how this contributes to indirect transmission of CWD.

### v2

@@ 1 Abstract
Chronic Wasting Disease (CWD) is a neurological disease impacting deer, elk, moose, and other cervid populations and is caused by a misfolded protein known as a prion. CWD is difficult to control due to the persistence of prions in the environment. Prions can remain infectious for more than a decade and have been found in soil as well as other environmental vectors, such as ticks and plants. Here, we provide a bifurcation analysis of a mathematical model of CWD spread in a cervid population, and use a modification of the Gillespie algorithm to explore if wolves can be used as an ecological control strategy to limit the spread of the disease in several relevant scenarios. We then analytically compute the probability that the disease spreads given one infected member enters a fully healthy population and the probability of elimination given a fully susceptible population and remaining prions in the environment. From our analysis, we conclude that wolves can be used as an effective control strategy to limit the spread of CWD in cervid populations, and hunting or other means of lowering the susceptible population are beneficial to controlling the spread of CWD, although it is important to note that inferring biologically relevant parameters from the existing data is an ongoing challenge for this system.
@@ 2 Introduction
Chronic Wasting Disease (CWD) is a prion-mediated neurodegenerative disease that has been observed in cervids, such as elk, deer, and moose. Prions are misfolded proteins that can cause neurological disease, such as CWD, but also Creutzfeldt-Jakob disease in humans. CWD is transmitted both directly between susceptible animals and indirectly through environmental exposure. Cervids can become infected through contact with saliva, blood, and other bodily fluids of an infected animal or through indirect environmental infections from water sources, plants, soil, or potentially vector-borne insect infections. CWD is known to influence the movement ecology of deer. A recent study showed that infected deer moved more slowly and were found in lower elevation areas compared with healthy deer that were not starving, and when compared to starving deer, the infected deer more often utilized areas near streams and with lower herbaceous biomass [6]. In recent years, wild populations of cervids have exhibited high levels of CWD prevalence. The 2024 Chronic Wasting Disease Surveillance Report from the Wyoming Game and Fish Department reported that the prevalence of CWD in some populations of mule deer is as high as 66.3%  [^1] The report can be found here https://wgfd.wyo.gov/wyoming-wildlife/wildlife-disease-and-health/chronic-wasting-disease using the drop-down Report menu at the bottom of the website.. CWD has proven difficult to control or eradicate, due to the time scale associated with the persistence of CWD prions in the environment. Prions decay slowly and can remain latent and infectious for more than a decade [12, 39]. Many mathematical models have been developed to help understand the disease spread in deer populations, evaluate the efficacy of control strategies, and infer the modes of transmission. We briefly review several of the published models, but our summary should not be considered exhaustive.
@@
One of the earlier discrete population models examined the spread of CWD and control strategies [25, 15], but the modeling assumptions of the models, in particular frequency-dependent transmission, were reevaluated in the follow-up paper [33]. Miller and co-authors developed and attempted parameter estimation of competing SIR-type models that describe transmission of CWD in a population of cervids through different mechanisms—though at least one of the models considered appears to exhibit a lack of structural identifiability, which we briefly comment on in the Discussion [24]. They found support for indirect transmission of CWD using a model selection approach [24]. A different study also suggests that the disease will spread via indirect transmission due to the persistence of prions in the environment [4]. The model presented in [24] was further modified for further mathematical and control-based investigations.
@@
Sharp and Pastor modified the model presented in [24] to include logistic growth in the susceptible population and performed a mathematical analysis of the system [34]. Wild and co-authors later extended the model presented in [24] to include the effects of predation from wolves [40] and analyzed the system in a non-oscillatory parameter regime, using parameter values estimated in [24] and other studies. An earlier report to the National Park Service by Hobbs also analyzed wolf-predation as a control strategy using a similar deterministic model in an oscillatory parameter regime [16]. Barbera and Pollino performed a mathematical analysis of a PDE [5] that was inspired by the model presented in [34]. Reyes and co-workers designed a spatiotemporal causal inference scheme to examine the effects of culling strategies using a PDE-based approach [22]. Potapov and co-authors used a complex age-structure differential equation model to examine the influence of harvesting strategies on population dynamics and disease suppression [30]. Vasilyva and co-authors performed a mathematical analysis of a differential equation model that describes CWD spread in a population of deer, including both indirect and direct transmission [37]. They also attempted a parameter estimation of the system, using upper and lower bounds for parameter estimation based on estimates of parameters found in [24]. A different study examined the influence of culling strategies to control CWD spread using a differential equation model that incorporated seasonality into a population model of cervids [29]. Others created an integrodifference equation model that incorporated direct and indirect transmission and long-distance dispersal from juvenile deer to understand CWD spread in a population of white-tailed deer [23]. Several groups have examined harvest strategies using mathematical models that delineate the deer population by sex [3, 2, 31]. Control investigations into CWD transmission in deer populations are summarized in [36].
@@ 4 Stochastic simulation
Here, $E$ remains a continuous, not discrete, variable, and so is governed by the differential equation $\frac{dE}{dt}=\frac{\epsilon A_{0}}{A}i-\mu_{e}E$. We would like to use these reaction rates with the Gillespie algorithm [14] to do stochastic simulations. However, since $E$ is continuously changing in time, the Gillespie algorithm is not directly applicable. Instead, we use “Poisson thinning” to account for this fact, as follows[20, 28, 32]: Since between reactions, $E(t)$ is a monotone function, bounded between $E(t_{0})$ and $\epsilon\frac{iA_{0}}{\mu_{e}A}$, where $t_{0}$ is the last reaction time, we let $E^{*}={\rm max}[E(t_{0}),\epsilon\frac{iA_{0}}{\mu_{e}A}]$, and introduce a modified reaction $R_{3}^{*}$ (since $R_{3}$ depends on $E$), with reaction rate $r_{3}^{*}=\gamma sE^{*}$ (which is the fastest possible reaction rate), and then proceed with the usual Gillespie determination of next reaction time and next reaction. However, whenever the next reaction is to be $R_{3}^{*}$ at time $t^{*}$, the reaction is split into two, only one of which is implemented: a null reaction for which no change to $s$ or $i$ is made with probability $\frac{E^{*}-E(t^{*})}{E^{*}}$, and the reaction that reduces $s$ by one and increases $i$ by one with probability $\frac{E(t^{*})}{E^{*}}$.
@@ 4.3 Probability of Elimination
is the probability that there will not be another outbreak. Obviously, the probability that there will not be another outbreak goes to zero as ${\gamma_{e}\over\mu_{e}}sE_{0}$ becomes large. This once again shows the risk associated with overpopulation of susceptible deer (large $s$), and the benefit of having wolves or other harvesting measures to help control the population size of healthy deer, even when there are no infected animals present. This is borne out by a recent statistical study that showed that deer populations that experienced high harvest pressure had lower prevalence of CWD [26].
@@ 5 Discussion
The prions responsible for CWD are known to exist in soil, and can be measured from soil samples [19]. In practice, estimating $E$ through soil samples would likely be expensive, infeasible on a large scale, and miss some of the sources of the prions in the environment. The prions responsible for CWD can be found in multiple environmental reservoirs beyond soil. It was recently shown that ticks can harbor levels of prions that could cause transmission of CWD [17]. Prions were also recently shown to exist in plants at relevant levels for transmission of CWD [7]. For these reasons, getting estimates of the prion density will be difficult. A different approach to ensure that unique parameters can be found is to reparameterize the model to remove the structural identifiability issue ahead of parameter estimation (for an example in a different modeling context, see [11]), but this can alter the interpretability of state variables of the model or rate parameters. Of course, whether unique parameter estimates are needed depends on how the model will be used.
@@
We chose to use an ODE model and did not consider the role of possible spatial inhomogeneity. However, the role of space may be important in understanding the spread of CWD in deer. Currently, little is known about the spatial distribution of prions. One would imagine the spatial distribution of the prion is somewhat “patchy” in space because an infected animal carcass contains a high concentration, which would then diffuse in the local environment. However, some degree of spatial mixing is likely to occur. For example, rivers could be an agent by which the prions could spread in the environment. Prions can remain infectious after passing through a crow digestive system [10, 38] and a coyote digestive system [27]. The same may be expected from a wolf, and so it may be expected that wolves would spread the prions throughout the environment through feces. In our model, we took the term $\xi W$ in Eq. 9 to be zero throughout our analysis, but taking this parameter to be non-zero would begin to capture prion spread by wolves. Currently, less is known about the spread of prions through insects and plants, or if infected plants can pass the prions inter-generationally through seeds, but these are other theoretical mechanisms by which prions could become well-mixed in the environment. More research is needed to understand the spatial distribution of prions in the environment and how this contributes to the indirect transmission of CWD.
@@ 6 Funding
7 Author Contribution
C.E.F. and J.P.K. designed the research. J.P.K. performed the model analysis and simulations. C.E.F. reviewed the analysis and performed the identifiability analysis. J.P.K. and C.E.F. wrote and edited the paper.
8 Code
The code used for this analysis can be found here https://github.com/cefitzg/cwd_code.git